### v2

# DepthART: Monocular Depth Estimation as Autoregressive Refinement Task

Abstract
Despite recent success in discriminative approaches in monocular depth estimation its quality remains limited by training datasets. Generative approaches mitigate this issue by leveraging strong priors derived from training on internet-scale datasets. Recent studies have demonstrated that large text-to-image diffusion models achieve state-of-the-art results in depth estimation when fine-tuned on small depth datasets. Concurrently, autoregressive generative approaches, such as the Visual AutoRegressive modeling (VAR), have shown promising results in conditioned image synthesis. Following the visual autoregressive modeling paradigm, we introduce the first autoregressive depth estimation model based on the visual autoregressive transformer. Our primary contribution is DepthART – a novel training method formulated as Depth Autoregressive Refinement Task. Unlike the original VAR training procedure, which employs static targets, our method utilizes a dynamic target formulation that enables model self-refinement and incorporates multi-modal guidance during training. Specifically, we use model predictions as inputs instead of ground truth token maps during training, framing the objective as residual minimization. Our experiments demonstrate that the proposed training approach significantly outperforms visual autoregressive modeling via next-scale prediction in the depth estimation task. The Visual Autoregressive Transformer trained with our approach on Hypersim achieves superior results on a set of unseen benchmarks compared to other generative and discriminative baselines.
Figure 1: We present the Depth Autoregressive Transformer for monocular depth estimation, trained using our novel procedure formulated as the Depth Autoregressive Refinement Task – DepthART. Our model iteratively enhances the depth map by predicting next-scale residuals, resulting in a highly detailed final estimate.
Introduction
Monocular depth estimation (MDE) is a fundamental problem in computer vision. Depth maps provide a compact intermediate scene representation useful for decision making in physical surroundings. Recovering depth data from a single image promises a high practical value for different applications including spatial vision intelligence (Wang et al. 2019; Godard et al. 2019), autonomous driving (Wang et al. 2019; Godard et al. 2019) and robotics (Wofk et al. 2019).
Early learning-based approaches (Eigen, Puhrsch, and Fergus 2014) tackle the monocular depth estimation problem as a supervised regression task. However, these methods were domain-specific (Silberman et al. 2012; Geiger, Lenz, and Urtasun 2012) and heavily relied on annotated datasets. As a result, they were subject to a limited generalization ability caused by a low amount of annotated data available. Recent techniques suggested different tricks challenging this limitation. MiDaS  (Ranftl et al. 2020) proposed to mitigate this issue by using an affine invariant depth training scheme on a mixture of datasets. While newer approaches proposing annotated data sources still appear (Yang et al. 2024), the acquisition of accurate depth annotations at scale remains challenging.
Recent studies (Ke et al. 2024; Fu et al. 2024) have highlighted the effectiveness of text-to-image diffusion models, originally trained on internet-scale image-caption datasets, as priors for monocular depth estimation. These approaches involve fine-tuning a pretrained diffusion model on a smaller, synthetic dataset with depth annotations, resulting in models that generate accurate and highly detailed depth maps. Concurrently, advancements in autoregressive models, such as the Visual AutoRegressive modeling (VAR) (Tian et al. 2024) and LLaMA-Gen (Sun et al. 2024), have demonstrated the capability of these models to generate high-quality images in class- or text-guided settings. These findings motivate an exploration of autoregressive generative techniques for depth estimation, offering a promising new direction.
In this work, we introduce a novel approach to monocular depth estimation based on the Visual AutoRegressive modeling (Tian et al. 2024). Our core contribution is the novel training procedure formulated as Depth Autoregressive Refinement Task. Our approach constructs dynamic targets using the model’s own predictions, rather than relying on ground truth token maps during training. By framing the objective as residual minimization and using model predictions as inputs, we bridge the gap between training and inference stages in autoregressive modeling, leading to enhanced depth estimation quality. We validate our model extensively comparing it with popular baselines under similar conditions. To the best of our knowledge, this is the first autoregressive depth estimation model. Moreover, it performs on-par or superior compared with popular depth estimation baselines.
Eventually, we formulate our contributions as follows:
  1. 1.
    We introduce a novel application of autoregressive image modeling for depth estimation by developing the depth autoregressive transformer.
  2. 2.
    We propose a new training paradigm for depth estimation, termed the Depth Autoregressive Refinement Task (DepthART), which facilitates self-refinement and incorporates multi-modal guidance during training.
  3. 3.
    We demonstrate, through extensive experiments, that the depth autoregressive transformer trained with DepthART achieves competitive or superior performance compared to existing baselines across several benchmarks not seen during training.
Related work
Monocular depth estimation
Learning-based monocular depth estimation approaches can be broadly categorized into two main branches: metric and relative depth estimation methods. Metric depth estimation (Laina et al. 2016; Alhashim and Wonka 2018; Bhat, Alhashim, and Wonka 2021, 2022; Yin et al. 2023) focuses on regressing absolute predictions at a metric scale. These models are typically trained on small, domain-specific datasets, which limits their ability to generalize efficiently across diverse environments. At the same time, relative depth estimation methods aim to estimate depth up to unknown shift and scale (SSI) or just unknown scale (SI). MiDaS  (Ranftl et al. 2020) introduced shift and scale invariant depth training on a mixture of several domain-specific datasets, significantly improving model generalization. Despite it, the depth predictions remained geometrically incomplete, i.e. point clouds cannot be built using model predictions. GP2  (Patakin et al. 2022) addressed this limitation by proposing an end-to-end training scheme that estimates a scale-invariant, geometry-preserving depth maps. Meanwhile, two-stage pipelines were developed to reduce shift ambiguity in depth maps at the second stage (Yin et al. 2021) or to upgrade the depth map to metric scale (Bhat et al. 2023). Further advancements in the field have been driven by the integration of various priors (Patil et al. 2022; Yang et al. 2024; Yin et al. 2023), improvements in architectural designs (Ranftl, Bochkovskiy, and Koltun 2021; Agarwal and Arora 2023; Ning and Gan 2023) and the expansion of training data (Yang et al. 2024).
Generative modeling
Recently, diffusion models have demonstrated their versatility across various computer vision tasks, including image generation  (Ho, Jain, and Abbeel 2020; Rombach et al. 2022; Ho et al. 2022), video generation  (Blattmann et al. 2023a; Bar-Tal et al. 2024; Blattmann et al. 2023b), or 3D objects modeling  (Poole et al. 2022; Wang et al. 2023b; Lin et al. 2023; Melas-Kyriazi et al. 2023). Beyond these applications, diffusion models have also been successfully employed in other problems, such as depth estimation (Saxena et al. 2023; Duan, Guo, and Zhu 2023; Saxena et al. 2024), image segmentation (Wang et al. 2023a; Amit et al. 2021) and object detection (Chen et al. 2023). Notably, Marigold  (Ke et al. 2024) and GeoWizard  (Fu et al. 2024) have demonstrated that the Stable Diffusion model (Rombach et al. 2022), pretrained on the large-scale image-caption dataset LAION-5B (Schuhmann et al. 2022), can produce high-quality depth maps after minor finetuning. This highlights the potential of utilizing pretrained generative models to enhance depth estimation accuracy and robustness across different domains.
Autoregressive modeling
While diffusion models remain one of the most widely-used generative approach, recent advancements in autoregressive models have shown significant promise for various generative tasks (Yu et al. 2023; Sun et al. 2024; Tian et al. 2024; Ma et al. 2024). These methods rely on discrete token-based image representations typically generated by VQ-VAE (Van Den Oord, Vinyals et al. 2017) or its derivatives. These derivatives often include architectural enhancements (Yu et al. 2021; Razavi, Van den Oord, and Vinyals 2019), additional masking techniques (Huang et al. 2023), or the incorporation of adversarial and perceptual losses (Esser, Rombach, and Ommer 2021). Autoregressive image synthesis is generally formulated as the sequential generation of tokens (Esser, Rombach, and Ommer 2021), followed by decoding them from the VQ space. Many approaches employ the GPT-2 (Radford et al. 2019) decoder-only architecture to predict sequences of VQ-VAE tokens (Esser, Rombach, and Ommer 2021; Chang et al. 2022; Ramesh et al. 2021; Chang et al. 2023; Yu et al. 2022). However recent works (Tian et al. 2024; Ma et al. 2024) have introduced the concept of predicting multi-scale token maps rather than token sequences. This approach reduces the risk of structural degradation and decreases the generation time for high-resolution images, enabling high-quality class- and text-conditioned image generation.
Preliminary
Figure 2: VQ-VAE (Tian et al. 2024) provides a single decomposition, while there are different trajectories resulting the same features. In contrast, DepthART enables multi-modal training, offering diverse refinement paths compared to the unimodal guidance used in the Visual AutoRegressive modeling (VAR).
Figure 3: We highlight the key differences between the original VAR approach (left) and our proposed training approach DepthART (right). In the VAR approach, quantized token maps provided by VQ-VAE serve as both inputs and targets during training. Our DepthART method introduces a refinement process (highlighted in the red box), where the model self-refines by using its predicted token maps as inputs instead of predefined VQ-VAE scales. The targets are defined as the quantized residuals between the encoded depth features $f_{D}$ and the cumulative model predictions up to the current scale. Depth features $f_{D}$ are extracted from the VQ-VAE encoder without undergoing quantization.
Next-scale visual autoregressive modeling.
Typically autoregressive image generation involves predicting image tokens in a raster scan order. However, recent work (Tian et al. 2024) introduced a novel autoregressive training approach for class-conditional image generation - visual autoregressive modeling. Instead of predicting tokens individually, they proposed generating token maps with varying scale. Each predicted token map progressively increases in resolution compared to the previous one, resulting in a scale-wise decomposition of the image. Specifically, an image ${\bm{I}}$ is modelled as a sequence of $\{x_{1},x_{2},\dots,x_{K}\}$, where $x_{k}$ is a token map of size $s_{k}=(h_{k},w_{k})$. The visual autoregressive transformer based on GPT-2 is then trained to maximize a class-conditioned likelihood:
$p(x_{1},x_{2},\dots,x_{K}|c)=\prod_{k=1}^{K}p(x_{k}|x_{1},x_{2},\dots,x_{k-1},c)$ (1)
Unlike the traditional token-by-token prediction approach, which can cause structural degradation due to raster scan ordering, visual autoregressive modeling predicts images scale-by-scale. Given that depth maps, like images, exhibit high spatial correlations, we explore the applicability of visual autoregressive modeling for solving depth estimation problem.
Discrete image representations.
Effective autoregressive image modeling relies on discretizing images into a finite set of tokens, a task closely related to image compression. Image compression methods have evolved from linear projections to advanced neural approaches, such as vector quantized variational autoencoders (VQ-VAE) (Van Den Oord, Vinyals et al. 2017). VQ-VAE approach combines an encoder-decoder neural network with a learnable token codebook of finite size. The encoder compresses the input image into a latent space, which is then quantized by mapping the latent vectors to the closest entries in the codebook. This quantized representation is then decoded back into the image space, with the entire process trained end-to-end using a log-likelihood objective over the image distribution.
Visual autoregressive model training requires multi-scale image decomposition represented as a sequence of token maps. Differently from original VQ-VAE (Van Den Oord, Vinyals et al. 2017) an input image ${\bm{I}}$ is quantized by encoder ${\mathcal{E}}$ into $K$ token maps $\{x_{1},x_{2},\dots,x_{K}\}$ with resolutions $\{(h_{k},w_{k})\}_{k=1,K}$. Quantization operation $Q[\cdot]$ is performed using the same codebook $Z$ regardless of scale. Combined using upsample-convolution operators $\eta_{i}$ token maps are assumed to sum up into continuous features ${\mathcal{E}}({\bm{I}})$. Accordingly, the $k$-th scale map is calculated as a scaled and quantized residual of the extracted features:
$r_{k}={\mathcal{E}}({\bm{I}})-\sum_{i=1}^{k-1}\eta_{i}(x_{i})$ (2)
$x_{k}={\mathcal{Q}}\big{[}{\mathcal{S}}(r_{k},s_{k})\big{]}$ (3)
Eventually, decoder ${\mathcal{D}}$ recovers reconstructed image $\hat{{\bm{I}}}$ from given token maps:
$\hat{{\bm{I}}}={\mathcal{D}}\bigg{(}\sum_{k=1}^{K}\eta_{k}(x_{k})\bigg{)}$ (4)
In our approach, we adapt a pretrained modification of VQ-VAE tailored specifically for the visual autoregressive modeling transformer (VAR) (Tian et al. 2024). While originally designed for colored images, we observe that VQ-VAE (Tian et al. 2024) can be applied to encode depth maps as well (see  fig. 1).
Method
In this work, we formulate the monocular depth estimation task as an image-conditioned autoregressive generation problem. Inspired by VAR, we develop the depth autoregressive transformer that generates depth token maps given image conditioning. As our primary contribution, we introduce the novel training procedure formulated as Depth Autoregressive Refinement Task – DepthART.
Figure 4: Depth autoregressive transformer trained with DepthART demonstrates superior reconstruction quality at every scale (left) compared to the standart VAR training procedure, achieving a 1.7x improvement in reconstruction quality. Since we do not finetune the VQ-VAE, its end-to-end reconstruction quality is shown as a soft limitation on reconstruction error. Additionally, the overall predicted token probability distributions exhibit higher average entropy (right), indicating that our training procedure promotes multi-modality.
Depth Autoregressive Refinement Task.
The original Visual AutoRegressive modeling (Tian et al. 2024) relies on scale-wise decomposition of image provided by pre-trained VQ-VAE encoder. During training the model predicts next-scale token map from the previous sequence of ground truth token maps. The guidance objective is a cross-entropy loss between VAR prediction and the same scale token map provided by VQ-VAE. Inputting ground-truth token maps results into discrepancies between training and inference process, and accumulation of errors during inference.
We address this issue and reformulate training objective as a Depth Autoregressive Refinement Task (DepthART). Our main goal is to enable model self-refinement during training. Hence, we construct inputs and targets dynamically from model predictions. Let’s consider an input image ${\bm{I}}$ with corresponding ground truth-depth map ${\bm{D}}$. We firstly encode an image into a series of token maps $\{x_{1},x_{2},\dots x_{K}\}$ provided by VQ-VAE (Tian et al. 2024). Resulting image token maps are fed to the model input as a starting sequence and serve as a conditioning for depth map estimation. Constructing dynamic supervision targets in our approach starts with performing model inference for given image token maps. We denote predicted depth token maps as $\{z_{1},z_{2},\dots,z_{K}\}$:
$z_{k}=\text{VAR}(z_{1},\dots,z_{k-1},x_{1},\dots x_{K})$ (5)
Next, we encode ground-truth depth ${\bm{D}}$ with the same VQ-VAE encoder into continuous features $f_{{\bm{D}}}$, discarding quantization process. Residual prediction targets $\{t_{1},t_{2},\dots,t_{K}\}$ can be constructed based on encoded depth features $f_{\bm{D}}$ and a series of models predictions up to current scale. This process is done similarly to VQ-VAE decomposition method (eqs. 2 and 3):
$\delta_{k}=f_{{\bm{D}}}-\sum_{i=1}^{k-1}\eta_{i}(z_{i})$ (6)
$t_{k}={\mathcal{Q}}\big{[}{\mathcal{S}}(\delta_{k},s_{k})\big{]}$ (7)
Eventually, the training objective takes the form of cross-entropy loss between predicted and target token maps:
${\mathcal{L}}=\sum_{k=1}^{K}{\mathcal{L}}_{CE}(z_{k},t_{k})$ (8)
As the result we form a new set of training samples, where $z_{k}$ and $t_{k}$ are model inputs and targets respectively. In contrast to original VAR training, these token maps are dynamically constructed at every training step rather than relying on a single predefined VQ decomposition (see fig. 3).
The training process benefits from such formulation in a few ways. Firstly, our procedure enables model self-refinement by making model aware of own predictions and framing the task as residual refinement. Since VQ-VAE (Tian et al. 2024) decomposition comes from approximating continuous features with summation of discrete token maps, multiple plausible decompositions can exist. Secondly, we argue that exploring various ways of decomposing input into token maps is beneficial for model training. The proposed training method eliminates single-mode limitation and facilities multi-modal solutions discovery (see fig. 2).
Depth Autoregressive Transformer
We develop our depth autoregressive transformer by leveraging the visual autoregressive transformer model, which is based on the GPT-2 architecture (Radford et al. 2019) and pretrained through visual autoregressive modeling (Tian et al. 2024) on the ImageNet dataset (Deng et al. 2009). Originally designed for class-conditioned image generation, we repurpose this model for image-conditioned depth map estimation. We emperically found that the optimal performance is achieved by inputting concatenated image tokens with previously predicted depth tokens. The complete training and inference pipeline is illustrated in fig. 3 (right).
[…]
GP-2 (EffNet-B5) 0.23 0.175 0.427 0.247 0.162 0.125 0.162 0.121 4.70 12.8 5.2
Midas (ResNeXt-101) 0.203 0.160 0.325 0.207 0.143 0.116 0.140 0.112 2.71 10.8 3.5
AdaBins (EffNet-B5) 0.235 0.184 0.323 0.206 0.141 0.115 0.161 0.125 4.65 12.6 4
DiT-depth (DiT) 0.309 0.220 0.252 0.169 0.149 0.120 0.169 0.127 2.86 8.71 4.3
DPT (ViT-L) 0.198 0.150 0.435 0.251 0.121 0.107 0.134 0.108 2.97 8.14 2.9
VAR (GPT-2) 0.245 0.285 0.396 0.294 0.185 0.141 0.177 0.133 1.98 9.44 6.1
DepthART (Ours, GPT-2) 0.196 0.177 0.275 0.178 0.141 0.115 0.129 0.106 1.91 7.27 2.1
Table 1: Quantitative evaluation across benchmarks not seen during training. Overall performance is summarized using a rank metric. Our depth autoregressive transformer, trained with DepthART, outperforms the original VAR training procedure and achieves the highest overall performance among a diverse set of depth estimation baselines.
Figure 5: Qualitative comparison of point clouds reconstructed from predicted depth maps on the IBIMS dataset. The depth autoregressive transformer trained with DepthART delivers higher-quality reconstructions, particularly in planar regions.
Figure 6: Qualitative comparison of the depth autoregressive transformer trained with DepthART against various baselines. Our model generates more precise depth estimates in planar regions while maintaining the overall scene structure.
Experiments
In this section, we empirically validate the effectiveness of our proposed approach on several depth estimation benchmarks that were not used during training. Specifically, we demonstrate that: (1) the DepthART training method significantly enhances the performance of the depth autoregressive transformer and introduces multimodality compared to the original VAR, (2) the depth autoregressive transformer trained with DepthART achieves comparable or superior accuracy relative to other baselines.
Implementation details
For our experimental evaluation, we chose to predict depth maps up to an unknown scale (SI). Unlike the commonly used scale-and-shift invariant training approach (Ranftl et al. 2020), scale-invariant training preserves the ability to reconstruct geometry from the predicted depth maps, which is essential for practical applications like single-view reconstruction. Therefore, prior to inputting the depth maps into VQ-VAE, we apply the following transformation:
${\bm{D}}_{norm}=\frac{{\bm{D}}}{{\bm{D}}_{98}+\epsilon}\times 2-1$ (9)
where $d_{98}$$98\%$ percentile of individual depth map.
Our depth autoregressive transformer is trained with DepthART using AdamW (Loshchilov and Hutter 2019) optimizer with a learning rate of $10^{-4}$ and weight decay of $10^{-2}$ and batch size equals to 4. Additionally we decrease learning rate during training with StepLR scheduler with a step size of $10,000$ and a gamma of $0.8$. Training of our model takes 2 days using 4 NVIDIA H100 GPUs.
Training protocol
To ensure consistent training conditions across all models, we train both the depth autoregressive transformer and baseline models on the same dataset. Due to the requirement of dense ground-truth depth maps for variational autoencoders, we utilize the highly realistic synthetic HyperSim dataset (Roberts et al. 2021), which includes 461 diverse indoor scenes. The pretrained VQ-VAE (Tian et al. 2024) used in our experiments generates multi-scale token maps only up to a maximum resolution of 256$\times$256, so we train all models at this resolution.
Evaluation protocol
Evaluation is performed on four datasets unseen during training: NYUv2 (Silberman et al. 2012) and IBIMS  (Koch et al. 2019) capturing indoor environments, TUM  (Li et al. 2019) capturing dynamic humans in indoor environment, ETH3D (Schops et al. 2017) providing high-quality depth maps for outdoor environments. Since all models trained to predict depth maps up to unknown scale, we first align predictions with ground-truth depth maps in terms of $\mathcal{L}_{1}$. Firstly, we evaluate accuracy of estimated depth maps using two commonly used metrics: Absolute Mean Relative Error (AbsRel) ($\downarrow$) and $\delta_{1}(\downarrow)$. Additionally, we assess the predicted depth maps using depth planar region deviations (pe-fla $\downarrow$) and plane orientation error (pe-ori, in ${}^{\circ},\downarrow$) on IBIMS dataset (Koch et al. 2019).
Baselines
We evaluate our approach against a diverse set of baseline models, organized into three categories. First, we consider several widely used depth estimation architectures trained discriminatively with an $\mathcal{L}_{2}$ regression loss, including MiDaS (Ranftl et al. 2020), GP2 (Patakin et al. 2022), and DPT (Ranftl, Bochkovskiy, and Koltun 2021). We also include AdaBins (Bhat, Alhashim, and Wonka 2021), which represents a classification-based approach to depth estimation. Additionally, we evaluate DiT (Peebles and Xie 2023), a transformer-based diffusion model pretrained on ImageNet. Finally, to assess the impact of our training procedure, we compare our depth autoregressive transformer trained with DepthART against the original VAR approach. More baseline training details are provided in the appendix.
Experimental results
VAR vs DepthART training.
To demonstrate the advantages of our approach, we trained our depth autoregressive transformer on the HyperSim dataset using both the original VAR training procedure and our DepthART method. Figure 4 presents a detailed comparison of these training methods, evaluated on the ETH3D dataset. We assess the reconstruction quality by calculating the AbsRel metric (fig. 4, left) between the intermediate depth maps (fig. 1, bottom) decoded at each autoregression step, based on cumulative predictions from both VAR and DepthART. Since we used a pretrained VQ-VAE without fine-tuning it for depth estimation, we also present its end-to-end depth map reconstruction error as a soft limit achievable by our depth autoregressive transformer. While the model trained using the VAR approach struggles to improve reconstruction quality at early scales, the DepthART-trained model consistently refines its predictions, achieving an overall reduction in relative error of approximately 70% compared to the baseline. Notably, the DepthART-trained model discovers a slightly better decomposition at the second scale than the reference provided by VQ-VAE, highlighting the non-uniqueness and potential suboptimality of the VQ-VAE decomposition. Additionally, we calculate the entropy of the predicted token distributions at each autoregression step (fig. 4, right). The higher entropy, coupled with the improved reconstruction quality, confirms that DepthART facilitates multi-modal training and leads to the discovery of more optimal prediction trajectories.
Comparison against baselines
To prove the efficiency of autoregressive approach in depth estimation, we train a set of popular baselines in similar conditions. Table 1 presents the evaluation results of all models on benchmarks that were not seen during training. To assess overall performance, we calculated each model’s rank on every dataset, and then averaged these ranks across all datasets. As can be seen from the Table 1, the depth autoregressive transformer trained with DepthART achieves the best overall performance. Notably, both models trained with the VAR and DepthART methods showed significantly better planar depth accuracy on the IBIMS dataset. We provide qualitative comparison of predicted depth maps in Figure 6. Besides, point clouds reconstructed from predicted depth maps (fig. 5) further support this observation. These results highlight the potential of autoregressive models for depth estimation.
Discussion
This work demonstrates the potential of generative autoregressive modeling for monocular depth estimation. Currently, our depth autoregressive transformer builds on VAR pretraining, which is constrained by the the ImageNet dataset. We believe that pretraining on more extensive and diverse datasets, such as those used for text-to-image generation, could significantly enhance our model’s performance. A primary limitation of our approach is the reliance on a VQ-VAE network derived from VAR, which has not been fine-tuned or retrained. This VQ-VAE was trained at low resolution on the relatively smaller OpenImages dataset (Kuznetsova et al. 2020), in contrast to larger, more recent datasets like LAION-5B (Schuhmann et al. 2022). We anticipate that upgrading to a higher-quality VQ-VAE could greatly benefit our method, and we identify these limitations as key directions for future research.
Conclusion
In this paper, we tackle the depth estimation problem through an autoregressive lens, specifically adapting the visual autoregressive modeling approach (Tian et al. 2024) for this task. Originally designed for class-conditioned image generation, we repurposed the visual autoregressive transformer for image-conditioned depth map estimation, introducing the Depth Autoregressive Transformer. Our analysis highlights limitations in the standard VAR training process, which leads to suboptimal accuracy on public depth benchmarks. To address these challenges, we proposed a novel training formulation, the Depth Autoregressive Refinement Task (DepthART). The Depth Autoregressive Transformer trained with DepthART showed substantial performance improvements over the VAR procedure and achieved competitive or superior results on public benchmarks compared to recent methods. Our approach enhances the model’s self-refinement ability and resolves the unimodality issues of visual autoregressive modeling, as demonstrated through empirical evaluation.
Appendix
Depth Autoregressive Transformer
Architecture
Originally, authors of VAR trained several models ranging from 300 million of parameters(depth=16) to 1 billion (width depth=24). We choose the smallest model with 300 million of parameters (depth=16) to make our model size compatible with other baselines. Secondly, since we train all models on a single synthethic dataset using lower number of parameters helps to avoid overfitting.
Inference
During the inference stage, we employed the top-k sampling algorithm to process the predicted token distribution. This method involves selecting the top-k token IDs from the distribution and subsequently sampling from this reduced set with replacement, generating outputs based on the multinomial distribution of the selected tokens. The use of top-k sampling allows for a more diverse range of generated tokens, which can help mitigate the deterministic nature of traditional methods such as argmax.
Baseline training
To evaluate the performance of our proposed model, we conducted a series of experiments involving multiple baseline models. All the models listed below were trained from ImageNet pretraining.
DPT, MiDaS, GP2
Three baseline models: the Dense Prediction Transformer (DPT) with a Vision Transformer (ViT) backbone with 344M parameters, the MiDaS model with a ResNeXt-101 backbone with 100M parameters, and the GP2 (EffNet-B5) model with 30M parameters, all optimized using $L_{2}$ loss on the Hypersim dataset at a resolution of $256\times 256$. For these models, we employed the AdamW optimizer, configuring it with a learning rate of $10^{-4}$ and a weight decay of 0.01. We also applied a StepLR scheduler with a step size of 10,000 iterations and a decay factor of 0.8 to manage the learning rate decay effectively.
DiT
A Diffusion Transformer (DiT) model with around 675M parameters, pretrained on ImageNet, was adapted for the task of depth prediction. To leverage the class-conditioned capabilities of the Diffusion Transformer, we concatenated the encoded depth and image latents along the feature dimension before passing them through the first convolutional projection layer. This modification effectively doubled the number of input channels, necessitating the duplication of the input layer’s weight tensor, with the weights halved to maintain the same initialization scale.
AdaBins
We use original model’s architecture with 78M parameters. We adhered to the original training methodology, incorporating the bin-center density loss proposed in the original work alongside the standard pixel-wise depth loss.
VQ-VAE
In this paper we used VQ-VAE derived from VAR paper. We found this model is able to encode depth maps successfully in most case. To encode it we replicate channel dimension to mimic RGB image and normalize depth value to a proper range. Still, We observe an instability issues of VQ-VAE. Specificially, this model can unpredictably corrupt encoding signal that contains errors in some pixels. Despite these values are in the valid range this model fails to reconstruct the underlying depth map from self encoded tokens. We consider these samples as out-of-distribution outliers since the VQ-VAE unlikely saw such differences in adjacent pixels during training. We demonstrate a couple of such outliers below.
Figure 7: By decoding depth maps on each autoregression step (after each sequentially predicted scale), we illustrate the depth refinement process introduced by our DepthART training problem formulation. Despite moderate improvmenets of numerical measures of reconstruction quality on finer scales (see fig.4 left in our paper), they are crucial for revealing precise object boundaries.
Figure 8: Qualitative comparison of point clouds reconstructed from depth predictions on IBIMS dataset
Figure 9: Qualitative comparison of point clouds reconstructed from depth predictions on NYU dataset
Figure 10: Qualitative comparison of depth maps generated by DepthART and baseline models on in-the-wild images from the COCOval2017 dataset. Despite all models were trained on synthetic indoor images and pre-trained on ImageNet, DepthART demonstrates superior generalization by producing more accurate depth predictions for out-of-distribution objects, including animals, cars, and people. Moreover, our model produces smoother depth gradients.